# Don’t Cut Corners: Exact Conditions for Modularity in Biologically Inspired Representations

### v1

@@ C.1 Structure of Representation - Act I
Where $F$ is smaller than half the number of neurons to ensure the autonomous linear system can propagate each frequency (each frequency ‘uses’ 2 dimensions, as will be obvious from Section C.3.
C.2 Readout Loss
@@
C.3 Recurrent Loss without Bias
@@
C.4 Recurrent Loss with Bias
@@
C.5 Structure of Representation - Act II - Large $\lambda_{W}$
@@ C.6 Frequencies
For any even, odd integer pair of $p,q\neq 1$ and $\delta>0$, $\cos(\pi/p)$ and $\cos(\pi/q)$ are bounded in $(0,1]$. In this case, the maximum possible value of the righthand side is achieved with the smallest possible pair of integers $(3,4)$, for which the inequality does not hold. Thus the inequalities eqn.132, 130 cannot hold at the same time which let us to conclude that there is no $\Delta$ below the critical value satisfying nonnegativity constraint and leads to modular representation.
The last case is where $\delta<0$ and both $p,q$ are odd. We can extend the relation shown in eqn.126 as following by simple substitution of oddity of the variables. With given odd integers $p,q$, there exists an even integer $k$ satisfying eqn.126. Similar to above, take $t=\frac{qk\pi}{p}$ for the mixed encoding eqn.123 and we develop:
@@
Appendix D Metrics for Representational Modularity and Source Statistical Interdependence
@@
Appendix E What-Where Task
E.1 Experimental Setup
E.1.1 Data Generation
@@ E.1.3 Training Protocols
E.2 Modularity of What & Where
E.2.1 Biological Constraints are necessary for Modularity
@@ E.3 Dropout & Correlation
Appendix F Nonlinear Autoencoders
F.1 Autoencoding Sources
@@ F.2 Autoencoding Images
Appendix G Recurrent Neural Networks
G.1 Linear and Nonlinear Periodic Wave RNN Experiment Details
@@
and used it as a proxy of mutual information for modularity metric introduced in eqn. 139.
G.2 Nonlinear Teacher-Student RNNs
@@
Appendix H Neural Data Analysis
We were kindly given the data of Xie et al. We used it to generate Figure 4B. In this section we talk through our analysis techniques. First, however we discuss methods for measuring subspace alignment.
H.1 Subspace Alignment Metrics
Previous works have calculated a single angle between subspaces, most rigorously the first principal angle [71], alternatively the unique angle between the two subspaces after projecting to a three dimensional PC space. We found that these angle based methods were biased for two reasons.
First, two $N$ dimensional subspaces have $N$ principal angles, and these are ordered, the first principal angle is smaller than the second, which is smaller than the third, etc. This ordering means that noise added to a representation will bias the estimates of the first principal angles down, and the last ones up. To show this we created data that lives in two planes oriented at 72 degrees. We then add guassian noise and find that this biases the estimate of the first principal angle down, and the second up, despite the fact that they are both 72, Figure 19.
Figure 19: We create some fake data of two 2D subspaces that are aligned at $72^{\circ}$, i.e. both of the two principal angles are $72^{\circ}$. We then add a small amount of noise 1000 times and calculate the principal angles between the noisy vectors (all the vectors are length one, and I add a zero-mean gaussian noise matrix with variance 0.1). As you can see, the estimate of the first principal angle is biased down, the second up, but the average appears unbiased.
Second, without further thought, angles are bounded, since they are necessarily smaller than 90. If your true data is oriented close to 90 degrees that means that noise will more likely push your estimate away from 90 than towards it. To show this effect we create another noisy dataset at 85 degrees, and show the noise pushes all estimates, even the mean of the two angles, significantly below 85, Figure 20.
Figure 20: Same setting as figure 19, except now the true principal angle is $85^{\circ}$. We can see that all estimators of the average principal angle are shifted downwards.
We counteract both these effects by using a single metric, whose bounds are nowhere near the ranges of interest. Specifically we use the correlation metric from Panichello & Buschmann. Given a set of vectors from the two subspaces of interest: $\{{\bm{g}}(c,s)\}_{c=1,s=1}^{C,2}$, indexed by subspace $s$ and condition $c$ (for example, which cue was presented), the correlation is:
$\rho=\langle\text{corr}({\bm{g}}(c,1)-\langle{\bm{g}}(c^{\prime},1)\rangle_{c^{\prime}},{\bm{g}}(c,2)-\langle{\bm{g}}(c^{\prime},2)\rangle_{c^{\prime}}\rangle_{c}$ (148)
Where the correlation is over neurons. This is 0 when subspaces are orthogonal, and avoids both pitfalls. Panichello & Buschman reported this metric on their data. We therefore move to extract the same metric from the data of Xie.
H.2 Extracting Coding Subspaces from Sampled without Replcement Sequences
[…]
Xie et al. get around this problem by doing regularised linear regression. To remove any potential hyperparameter dependence we employ a novel difference scheme. Consider the following term:
$\Big{\langle}{\bm{g}}\Big{(}\begin{bmatrix}\theta_{1}\\ \theta_{2}\end{bmatrix}\Big{)}-{\bm{g}}\Big{(}\begin{bmatrix}\theta^{\prime}_{1}\\ \theta_{2}\end{bmatrix}\Big{)}\Big{\rangle}_{\theta_{2}}$ (149)
Equal to:
$\frac{Q-2}{Q}\Big{\langle}{\bm{g}}\Big{(}\begin{bmatrix}\theta_{1}\\ \theta_{2}\end{bmatrix}\Big{)}-{\bm{g}}\Big{(}\begin{bmatrix}\theta^{\prime}_{1}\\ \theta_{2}\end{bmatrix}\Big{)}\Big{\rangle}_{\theta_{2}\neq\theta_{1},\theta^{\prime}_{1}}+\frac{1}{Q}\Big{(}\underbrace{{\bm{g}}\Big{(}\begin{bmatrix}\theta_{1}\\ \theta_{1}\end{bmatrix}\Big{)}-{\bm{g}}\Big{(}\begin{bmatrix}\theta^{\prime}_{1}\\ \theta_{1}\end{bmatrix}\Big{)}+{\bm{g}}\Big{(}\begin{bmatrix}\theta_{1}\\ \theta^{\prime}_{1}\end{bmatrix}\Big{)}-{\bm{g}}\Big{(}\begin{bmatrix}\theta^{\prime}_{1}\\ \theta^{\prime}_{1}\end{bmatrix}\Big{)}}_{\text{Error Term}}\Big{)}$ (150)
Again, the left hand side is what we want, the first term on the right is something we can get, and the leftover part is going to be ignored, our error term. Why have we chosen to leave this particular term out? Using the decomposition property of our representations we can see that the thing we want to estimate is:
$\Big{\langle}{\bm{g}}\Big{(}\begin{bmatrix}\theta_{1}\\ \theta_{2}\end{bmatrix}\Big{)}-{\bm{g}}\Big{(}\begin{bmatrix}\theta^{\prime}_{1}\\ \theta_{2}\end{bmatrix}\Big{)}\Big{\rangle}_{\theta_{2}}={\bm{g}}_{1}(\theta_{1})-{\bm{g}}_{1}(\theta^{\prime}_{1})$ (151)
Which makes sense, it is the difference in subspace 1 between the encoding of these two stimuli. But now we can analyse the error term and see that it is equal to the same thing!
${\bm{g}}\Big{(}\begin{bmatrix}\theta_{1}\\ \theta_{1}\end{bmatrix}\Big{)}-{\bm{g}}\Big{(}\begin{bmatrix}\theta^{\prime}_{1}\\ \theta_{1}\end{bmatrix}\Big{)}+{\bm{g}}\Big{(}\begin{bmatrix}\theta_{1}\\ \theta^{\prime}_{1}\end{bmatrix}\Big{)}-{\bm{g}}\Big{(}\begin{bmatrix}\theta^{\prime}_{1}\\ \theta^{\prime}_{1}\end{bmatrix}\Big{)}=2({\bm{g}}_{1}(\theta_{1})-{\bm{g}}_{1}(\theta^{\prime}_{1}))$ (152)
Hence, plugging these into equation 149:
$\Big{\langle}{\bm{g}}\Big{(}\begin{bmatrix}\theta_{1}\\ \theta_{2}\end{bmatrix}\Big{)}-{\bm{g}}\Big{(}\begin{bmatrix}\theta^{\prime}_{1}\\ \theta_{2}\end{bmatrix}\Big{)}\Big{\rangle}_{\theta_{2}\neq\theta_{1},\theta^{\prime}_{1}}={\bm{g}}_{1}(\theta_{1})-{\bm{g}}_{1}(\theta^{\prime}_{1})$ (153)
So we can see that this object is a good estimator of this quantity!
This is therefore the quantity we estimate. For all different stimuli pairs $\theta,\theta^{\prime}$, we build an estimate of the difference vectors. We then use these vectors to calculate the correlation alignment measure 148, this result is reported in Figure 4B. We form error bars by subsampling by half the number of neurons 200 times and estimating the mean and variance over 200 simulations.
H.3 Extracting Encodings from Single-Element ‘Sequenece’ Data
Finally, we need to choose how to encode the memories in the Xie task for our RNN models to use, see Appendix I. We take a data driven approach and estimate the monkey’s internal representation of each cue by analysing the trials in which the only a single stimulus was presented. We average these trials to create six encoding vectors, then find the dot product similarity matrix of this data. Finally, we find a set of 5 dimensional encodings that exactly recreate this data dot product structure. These are the vectors we feed our RNN: they have the same similarity structure as the monkey’s representation, without increasing training time by introducing a neuron-dimensional weight matrix.
Appendix I RNN Models of Neuroscience Tasks
We now talk through the biologically inspired linear RNNs used in sections 5 onwards of the main paper.
I.1 Xie Task, Model, and Extended Results
Each RNN is trained on many sequences. For each sequence the RNN sees two cue encodings, sampled from a set of six $\{{\bm{e}}_{i}\}_{i=1}^{6}$. During the first two timesteps the model is shown the two cues, then recieves no further input. Then there is a delay timestep. Finally in the last two timesteps the model must use an affine output to recreate the two encodings. The recurrent dynamics are simple linear systems:
${\bm{g}}_{t+1}={\bm{W}}_{\text{rec}}{\bm{g}}_{t}+{\bm{W}}_{\text{in}}{\bm{e}}_{t,i_{t}}+{\bm{b}}\qquad{\bm{e}}_{t,i_{t-3}}={\bm{W}}_{\text{out}}{\bm{g}}_{t}+{\bm{b}}_{\text{out}}$ (154)
${\bm{g}}_{0}=0$. We train the networks using gradient descent to perfectly solve the task while ensuring the representation is positive. Further, we penalise the L2 norm of the activity and the weights. We used a single set of hyperparameters to generate all data that can be found in the supplementary code.
We played with three different memory encoding schemes, either a 6 dimensional one-hot code, 6 two-dimensional points drawn evenly from the unit circle (like the positions of the dots on the screen shown to the monkey), or the data driven encoding discussed in the Appendix H. Finally, we used two different sequence structures: sampling both sequence elements uniformly, or sampling them without replacement but otherwise uniformly, as in the experiments.
We analysed the resulting delay period activity using the techniques discussed in Appendix H. We found, first, that in all cases, the delay period activity for networks trained on sequences sampled with replacement comprised two orthogonal subspaces. Second, depending on the encoding scheme, some networks trained on sequences sampled without replacement were orthogonal, others aligned, and the degree of alignment depending on the hyperparameter weighting of the activity vs. weight loss. Third, in both sequence settings we also found that the subspace encoding the first memory was a small amount bigger than the second, matching data. We now talk through some of these results in more detail.
Figure 21: We estimate the two-dimensional subspaces encoding, the two subspaces are basically orthogonal, the circular mean of the 2 principal angles between the subspaces is $89.95^{\circ}$
First, of the three encoding schemes we find that the 2D data is orthogonalised when sampled without replacement (Fig 21), while the one-hot (not shown) and data driven codes do not (Figure 4B). It was this puzzling discrepancy between the behaviour of the one-hot and 2D codes that led us to the data-driven encoding. This behaviour can be somewhat understood by looking at the joint distribution of the encoded memories, figure 22. The required inequality constraint is not satisfied, while it is for 1-hot memories. The neural data inspired encoding presumably lies closer to the 1-hot end of the spectrum.
Figure 22: Left: We plot the joint distribution of one dimension of each of the memories against one another for the 2D coding. The x axis is cosine of the first sequence element, the y axis is the second. Sampling without replacement removes the two extreme corner points (but not the whole diagonal, as though this dimension of the code might be equal, the memory is 2D, and the other dimension can be different). The line marks one of the inequality constraints that can be broken to induce mixing. As can be seen, removing one data point is not sufficient to remove all data below this line, so the code modularises. Right: The converse is true for one-hot variables.
We return to the alignment of correlated sequences after discussing the similar results from the Panichello & Buschmann studies.
I.2 Panichello & Buschmann, Model, and Extended Results
We study a very similar linear RNN setting, the loss, training protocol, and sequence structures are mostly identical to the previous section. We list the differences. First, we assume the memory encoding is the equivalent of the 2D one from the previous section.
Second, network architecture: rather than providing inputs to the linear network, we let the network learn the input by allowing it to initialise itself differently depending on the pair of stimuli. This was an implementational choice that doesn’t constrain the input to be a linear function of the encodings. Nonetheless, the network learns to initialise itself as a linear function of the two presented cues. We could equally have provided the two cues as inputs as in the previous section, it would likely have changed little.
Third, we implement the different cues by updating the initial representation with different matrices. If the animal is cued to attend to the top stimulus the initial activity, which we think of as the delay period activity, ${\bm{g}}_{0}(c_{\text{top}},c_{\text{bottom}})$ for two colours $c_{\text{top}}$ and $c_{\text{bottom}}$, is updated using one matrix, which then drives the readout:
${\bm{W}}_{\text{top}}{\bm{g}}_{0}(c_{\text{top}},c_{\text{bottom}})+{\bm{b}}_{\text{top}}={\bm{g}}_{1}(c_{1},c_{2})\quad{\bm{R}}{\bm{g}}_{1}(c_{\text{top}},c_{\text{bottom}})+{\bm{b}}_{\text{out}}={\bm{e}}_{c_{\text{top}}}$ (155)
And the same for the bottom matrix. This is inspired by recent work that has shown these type of action dependent matrices are both mathematically tractable, and biologically plausible [40, 14, 68, 70, 69].
Fourth, Number of Samples: in the real experiment Panichello & Buschman present colours drawn from a continuous colour wheel. They then analyse it by binning the colour wheel into groups of four. We skip straight to the binned colours, and pretend the encoded cues that the network predicts are 4 points sampled evenly from the unit circle. Other than this change the sampling schemes are the same.
We extract the subspaces and their alignemnts as discussed in Appendix H. We make some final interesting (to us!) observations. First, the top and bottom colour subspaces are equally sized, unlike in Xie et al., likely because they are equally important, whereas in the Xie task one stimuli has to be recalled before the other, making it, temporally, more important. Second, since we only sample 4 ‘colours’ from the circle, rather than the 6 in Xie et al., the prediction shown in figure 22 changes. With four data points even 2D encodings are missing the appropriate corner 23!
Figure 23: As in 22, we plot the joint distribution of one dimension of each of the memories against one another for the 2D coding. The x axis is cosine of the first sequence element, the y axis is the second. Since there are now only 4 settings of the angle we can see sufficient corner has been removed to predict mixing, as we see in simulations Figure 4.
This matches the finding presented in Figure 4, that when sampled without replacement the subspaces align, since a sufficiently large corner is missing.
Finally, there remain discrepancies between the neural data and our theory and models. Most fundamentally, both theory and models predict that if the subspaces are orthogonal they should be encoded by disjoint groups of neurons, if they align, they should not. Yet, both Xie and Panichello and Buschman find neurons tuned to both cues. This is expected for Xie’s data, but not for Panichello and Buschman’s, where they find roughly 20% of colour tuned neurons are mixed selective. We do not yet have a good understanding of this discrepancy. More superificially, we are not able to fit the exact numerical value of the alignment between subspaces in Xie et al.’s data, and we remain unsure whether this could be done with different hyperparameters or a noise model. We leave this for further exploration in a more targeted venue.
I.3 Motor Subspace Model
To study preparatory and potent subspaces we train ReLU RNNs on simple reach tasks. Each reach is defined as a one-hot code, such as a location to reach towards. The one-hot part of the network’s input tells it which reach it will have to perform next, but not when. A final dimension of the input encodes the go signal saying ’on the next time step do the cued reach!’. The network is then trained to output the cued-one-hot code for a couple of timesteps after the go cue. The network is either trained on a variety of single-reaches, or two reaches cued one after the other.
We then measure how aligned the preparatory and potent subspaces are. Following [21, 75] we gather the activity in the network during single-reach tasks and break it into preparatory - activity after the cue has been presented but before the network outputs anything - and potent - activity during the reach - activity. We then perform PCA on these two sets of activity. We build subspaces made from a subset of these principal components: if the neural activity projected along a principal component captures more than $0.1\%$ of the total variance of that neural activity (preparatory or potent) we include it. We then measure how much of one type of neural activity (preparatory or potent) is captured after projecting to the other subspace.
I.4 Warden & Miller Task, and Model
In the task two images sampled without replacement from a set of four are shown to the animal with delay periods in between. The animal then has to report its memory of the images in a couple of different ways.
We use the same linear RNN as before, but with three differences. First, we include a delay period between the first and second cue, and the second cue and the report time. Second, to take into account the heterogenous nature of the report, we simply ask the network to output 2 4-dimensional one-hot vectors, corresponding to the images presented first and second. Third, we use a different weight update matrix during delay times and cue presentation times. So at times C1 and C2 (Figure 7 H, I) we use one recurrent weight and bias, during the delay period we use another. This was to enable such a simple linear system to solve the task in a minimal way.
To create the plots we perform the same analysis as [41], we average the neural firing rates at each timepoint according to which stimulus was presented first or second. We find many neurons that code for stimulus-time conjunctions, and fewer that code for only time, and they tended to fire around cue times. Note these averages do not take account of the correlations between the two images, introduced by sampling without replacement. As such, the tuning curves cannot be interpreted as evidence of mixed selectivity: in fact these neurons are modular, they respond only to a single stimulus. The other bumps in firing are induced by the task structure. Since the sequences are sampled with replacement our networks have other neurons that are linearly mixed selective to the two cues, as you would expect from our theory.
I.5 Entorhinal Model for Missing Variables Task, and Mixed Spatial-Reward Coding
Our final biological RNN model is inspired by the entorhinal literature and linear network models of it [14, 68, 70, 69]. We used this model to model mixed spatial and reward coding Section 7, and mixed selectivity induced by missing correlated objectsSection 8.
@@ I.5 Entorhinal Model for Missing Variables Task, and Mixed Spatial-Reward Coding
In each trial the agent navigates a 3x3 periodic environment. The RNN receives one special input only at timepoint 0 that tells it the layout of the room. Otherwise at each timestep it is told which action (north, south, east, west), the agent took. It uses this action to linearly update its hidden state:
@@
In the spatial-reward task, at timepoint 0 the agent starts at a consistent position and receives an input that is a 9-dimensional 1-hot code telling the agent the relative position of the reward. At each timestep the agent then has to output two 9-dimensional 1-hot codes, one signalling its position in the room, the other the relative displacement of the reward from its position.
@@
In the mixed-selectivity task there are three objects in each room and the agent has to report the relative displacement of each of them. The input at the start of each trial is a 27-dimensional code signalling where the three objects are relative to its own position. As such, as the agent moves around the room it has to keep track of the three objects. This task was harder to train so we moved from the mean squared error to the cross-entropy loss and found it worked well, apart from that all details of loss and training are the same.
@@
We measured the linear NCMI between the neural activity and the 27-dimensional output code and found that each neuron was informative about a single source, as expected, it had modularised Figure 7C. However, pretend we did not know the third object existed. Perhaps it is a location the mouse cares about, but no experimenter could ever be expected to guess, such as its favourite corner, toilet spot, or scratching post - all good things to keep oriented in your mind. Instead we would calculate the linear NCMI between each neuron and those objects we know to exist. We would still find modular codes for these objects, but we would also find that the neurons that in reality code for third object, due to the correlations between object placements, are actually informative about the first and second object, so they look mixed selective! The position of the objects is always informative about each other, but by conditioning on each object we are able to remove this effect with our metric and uncover the latent modularity. But without knowing which latents to condition on we cannot proceed, and instead get lost in correlations.

### v2

@@ E.1 Structure of Representation - Act I
Where $F$ is smaller than half the number of neurons to ensure the autonomous linear system can propagate each frequency (each frequency ‘uses’ 2 dimensions, as will be obvious from Section E.3.
E.2 Readout Loss
@@
E.3 Recurrent Loss without Bias
@@
E.4 Recurrent Loss with Bias
@@
E.5 Structure of Representation - Act II - Large $\lambda_{W}$
@@ E.6 Frequencies
For any even, odd integer pair of $p,q\neq 1$ and $\delta>0$, $\cos(\pi/p)$ and $\cos(\pi/q)$ are bounded in $(0,1]$. In this case, the maximum possible value of the righthand side is achieved with the smallest possible pair of integers $(3,4)$, for which the inequality does not hold. Thus the inequalities eqn.144, 142 cannot hold at the same time which let us to conclude that there is no $\Delta$ below the critical value satisfying nonnegativity constraint and leads to modular representation.
The last case is where $\delta<0$ and both $p,q$ are odd. We can extend the relation shown in eqn.138 as following by simple substitution of oddity of the variables. With given odd integers $p,q$, there exists an even integer $k$ satisfying eqn.138. Similar to above, take $t=\frac{qk\pi}{p}$ for the mixed encoding eqn.135 and we develop:
@@
Appendix F Metrics for Representational Modularity and Source Statistical Interdependence
@@
Appendix G What-Where Task
G.1 Experimental Setup
G.1.1 Data Generation
@@ G.1.3 Training Protocols
G.2 Modularity of What & Where
G.2.1 Biological Constraints are necessary for Modularity
@@ G.3 Dropout & Correlation
Appendix H Nonlinear Autoencoders
H.1 Autoencoding Sources
@@ H.2 Autoencoding Images
Appendix I Recurrent Neural Networks
I.1 Linear and Nonlinear Periodic Wave RNN Experiment Details
@@
and used it as a proxy of mutual information for modularity metric introduced in eqn. 151.
I.2 Nonlinear Teacher-Student RNNs
@@
Appendix J RNN Models of Entorhinal Cortex
@@
We now talk through the linear RNNs used in Section 5. Our model is inspired by the entorhinal literature and linear network models of it [11, 55, 57, 56]. In each trial the agent navigates a 3x3 periodic environment. The RNN receives one special input only at timepoint 0 that tells it the layout of the room (format described later). Otherwise at each timestep it is told which action (north, south, east, west), the agent took. It uses this action to linearly update its hidden state:
@@
In the spatial-reward task, Figure 4, at timepoint 0 the agent starts at a consistent position and receives an input that is a 9-dimensional 1-hot code telling the agent the relative position of the reward. At each timestep the agent then has to output two 9-dimensional 1-hot codes, one signalling its position in the room, the other the relative displacement of the reward from its position.
@@
In the mixed-selectivity task, Figure 5, there are three objects in each room and the agent has to report the relative displacement of each of them. The input at the start of each trial is a 27-dimensional code signalling where the three objects are relative to its own position. As such, as the agent moves around the room it has to keep track of the three objects. This task was harder to train so we moved from the mean squared error to the cross-entropy loss and found it worked well, apart from that all details of loss and training are the same.
@@
We measured the linear NCMI between the neural activity and the 27-dimensional output code and found that each neuron was informative about a single source, as expected, it had modularised Figure 5C. However, pretend we did not know the third object existed. Instead we would calculate the linear NCMI between each neuron and those objects we know to exist. We would still find modular codes for these objects, but we would also find that the neurons that in reality code for third object, due to the correlations between object placements, are actually informative about the first and second object, so they look mixed selective! The position of the objects is always informative about each other, but by conditioning on each object we are able to remove this effect with our metric and uncover the latent modularity. But without knowing which latents to condition on we cannot proceed, and instead get lost in correlations.